# Network Analytics for Anti-Money Laundering – A Systematic Literature Review and Experimental Evaluation

### v2

@@ 5 Results and Discussion
Intrinsic features $0.9621\pm 0.0185$ $0.1702\pm 0.0035$ $0.2892\pm 0.0059$
Egonet features $0.9155\pm 0.0273$ $0.1620\pm 0.0049$ $0.2752\pm 0.0083$
Deepwalk-NI $0.0426\pm 0.0198$ $0.0075\pm 0.0035$ $0.0128\pm 0.0060$
Deepwalk $0.9733\pm 0.0159$ $0.1721\pm 0.0031$ $0.2925\pm 0.0051$
Node2vec-NI $0.0159\pm 0.0122$ $0.0028\pm 0.0022$ $0.0048\pm 0.0037$
Node2vec $0.9694\pm 0.0166$ $0.1715\pm 0.0032$ $0.2914\pm 0.0053$
GCN $\boldsymbol{1.0000\pm 0.0007}$ $\boldsymbol{0.2213\pm 0.0162}$ $\boldsymbol{0.3621\pm 0.0217}$
GraphSAGE $\boldsymbol{1.0000\pm 0.0000}$ $0.1765\pm 0.0007$ $0.3001\pm 0.0010$
GAT $0.9693\pm 0.0171$ $0.1711\pm 0.0073$ $0.2909\pm 0.0108$
GIN $0.9895\pm 0.0098$ $0.1746\pm 0.0069$ $0.2968\pm 0.0100$
@@
For the threshold-independent metrics, we focus on the AUC-PR, since this is more appropriate for high class imbalance. The models with intrinsic and those with egonet features give already good results. The suspicious transactions can be detected quite well using these “simpler” features. Nevertheless, the more advanced GNNs seem to bring additional predictive power. We see that GraphSAGE performs best, with GCN and GAT the second and third best models, respectively. For the threshold-depending metrics, the classic GCN is performing best on the smaller percentages, while GraphSAGE is best for the larger percentages. The additional value for GNNs in terms of precision, recall and F1 is much higher. This additional discriminatory power is important when resources to investigate transactions are limited. The relatively poor performance of GIN compared to the other GNNs is surprising.
In Table 10, we have added the results from models trained only on the embedding from the shallow representation learners, indicated with “NI” (no intrinsic features). We observe a significant performance drop in both AUC-ROC and AUC-PR due to the transductive nature of shallow representation methods. In the Elliptic data set, each time step features a distinct network, resulting in disparate embeddings. These embeddings represent distinct latent space coordinates, rendering them unsuitable for consistent insights across different parts of the data. Thus, we conclude that shallow representation is inadequate for addressing the dynamic challenge of detecting money laundering.
@@ 6 Conclusion
In this paper, we present the results of a systematic literature review of the literature on network analytics for anti-money laundering and of an experimental evaluation of the state-of-the-art. The systematic literature review resulted in a selection of 97 papers within-scope. To gain both comprehensive and deep insight, we analysed the global characteristics of these 97 papers, and carried out a more detailed analysis as well of the top-$10\%$ cited papers.
@@
We find that a large part of the literature concerns methods that are specifically tailored to solving just a small part of the money laundering puzzle. This might indicate a latent consensus that one should not aim at developing a single, overarching model for detecting money laundering. Instead, one may benefit from operating multiple and complementary models in parallel, each detecting different money laundering patterns. Additionally, most methods find patterns that are indicative for money laundering by relying on centrality metrics and other manually engineered features. The choice for these methods seems mostly motivated by the lack of interpretability of more complex network representation methods.
To extend upon the existing body of knowledge, we implemented an extensive experimental evaluation to compare the methods as proposed in literature. These were complemented by testing a series of alternative approaches. From the results, we can conclude that, although manually constructed network features give fairly good results, GNNs generally perform best. In terms of performance, a trade-off is to be made between predictive power and interpretability. Whereas manually engineered features are highly interpretable, they are time-consuming to conceive, implement and maintain. Alternatively, GNNs automatically extract predictive patterns from the data, but these patterns are hard to extract and interpret from the resulting black box models.
Based on the results of our work, we suggest the following topics for future research. For unsupervised learning, future research can focus on building a framework to compare different pattern matching methods. From the literature, it is also apparent that a comparative study for unsupervised AML methods is missing. Additionally, more research efforts could be invested in deep learning methods for unsupervised AML.
@@
For network visualisation, future work should focus on quantifying the usefulness of visualisations in helping investigators detect new patterns and money laundering cases. This can be extended into a comparative study that looks at the strengths and weaknesses of different visualisations.
Finally, in line with the conclusions of [69], work on the interpretation of GNNs in AML is still lacking. Future work should focus on applying known and novel methods on money laundering networks. This can help in uncovering previously unknown network structures used to launder money.
CRediT authorship contribution statement
Bruno Deprez: Conceptualization, Methodology, Data curation, Software, Formal analysis, Writing – original draft. Toon Vanderschueren: Conceptualization, Writing – review & editing. Wouter Verbeke: Conceptualization, Writing – review & editing. Bart Baesens: Conceptualization, Writing – review & editing. Tim Verdonck: Conceptualization, Writing – review & editing.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgments
This work was supported by the Research Foundation – Flanders (FWO research project 1SHEN24N) and by the BNP Paribas Fortis Chair in Fraud Analytics.
@@ Appendix A Literature Overview
Table 1: Literature classification. Classification of the papers according to different categories: publication data, learning method, evaluation metric, objective, and data.
Publication Data Learning Method Evaluation Metric Objective Data
@@
Reference Year Review Paper Bitcoin/Crypto Supervised Unsupervised Semi-Supervised Mixed Visualisation Accuracy Precision Recall F1 TPR FPR AUC-ROC AUC-PR Time Client Classification Transaction Classification Community Detection Flow/Chain Detection Proprietary Open-Source Synthetic
[31] 2015
[91] 2011
[79] 2013
[38] 2017
[139] 2017
[27] 2011
[80] 2016
[105] 1995
[125] 2021
[39] 2007
[63] 2018
[59] 2018
[85] 2011
[126] 2021
[72] 2020
[97] 2018
[142] 2017
[29] 2019
[64] 2019
[21] 2017
[119] 2021
[89] 2016
[138] 2014
[58] 2019
[114] 2019
[93] 2012
[76] 2015
[55] 2016
[133] 2021
[98] 2018
[41] 2021
[82] 2013
[128] 2021
[69] 2021
[78] 2021
[107] 2021
[115] 2020
[101] 2020
[13] 2018
[25] 2017
[140] 2022
[43] 2018
[77] 2018
[112] 2021
[62] 2018
[83] 2014
[129] 2018
[87] 2022
[6] 2020
[108] 2021
[99] 2017
[137] 2021
[42] 2022
[54] 2021
[84] 2019
[15] 2019
[92] 2014
[124] 2020
[110] 2021
[51] 2017
[49] 2021
[4] 2019
[23] 2007
[68] 2020
[60] 2022
[14] 2021
[109] 2022
[19] 2022
[61] 2022
[67] 2015
[111] 2022
[44] 2022
[32] 2022
[136] 2022
[40] 2021
[121] 2022
[45] 2022
[73] 2022
[71] 2022
[10] 2022
[104] 2017
[28] 2012
[22] 2022
[26] 2021
[7] 2020
[11] 2018
[66] 2011
[75] 2022
[100] 2022
[74] 2012
[131] 2020
[106] 2018
[134] 2022
[130] 2021
[81] 2013
[24] 2010
[53] 2021

### v3

@@ 5.1 Supervised Learning
Intrinsic features $0.874\pm 0.032$ $0.155\pm 0.006$ $0.263\pm 0.010$
Egonet features $0.928\pm 0.025$ $0.164\pm 0.005$ $0.279\pm 0.008$
DeepWalk $0.870\pm 0.034$ $0.154\pm 0.006$ $0.262\pm 0.010$
Node2vec $0.973\pm 0.016$ $0.172\pm 0.003$ $0.292\pm 0.005$
GCN $\boldsymbol{1.000\pm 0.000}$ $\boldsymbol{0.177\pm 0.007}$ $\boldsymbol{0.301\pm 0.010}$
GraphSAGE $0.981\pm 0.014$ $0.174\pm 0.006$ $0.295\pm 0.009$
GAT $\boldsymbol{1.000\pm 0.000}$ $0.176\pm 0.006$ $0.299\pm 0.009$
GIN $0.977\pm 0.019$ $0.172\pm 0.006$ $0.293\pm 0.009$
Table 14: Threshold-dependent metrics: Precision, recall and F1-score values for the top $1\%$ scores over the different methods for the IBM-AML data set, based on the test set. The standard deviation is also reported.
IBM-AML Supervised
Methods Precision Recall F1-score
Intrinsic features $0.021\pm 0.005$ $0.024\pm 0.005$ $0.022\pm 0.005$
Egonet features $0.020\pm 0.005$ $0.023\pm 0.005$ $0.021\pm 0.00$
DeepWalk $0.0209\pm 0.0044$ $0.0231\pm 0.0049$ $0.0220\pm 0.0046$
Node2vec $\boldsymbol{0.0214\pm 0.0050}$ $0.0237\pm 0.0055$ $\boldsymbol{0.0225\pm 0.0052}$
GCN $0.0010\pm 0.0010$ $0.0012\pm 0.0011$ $0.0011\pm 0.0011$
GraphSAGE $0.0094\pm 0.0003$ $\boldsymbol{1.0000\pm 0.0000}$ $0.0186\pm 0.0006$
GAT $0.0011\pm 0.0011$ $0.0012\pm 0.0012$ $0.0012\pm 0.0011$
GIN $0.0210\pm 0.0040$ $0.0231\pm 0.0042$ $0.0220\pm 0.0041$
For the threshold-independent metrics, we focus on the AUC-PR, since this is more appropriate for high class imbalance. For both data sets, the models with intrinsic and those with egonet features give already good results. The suspicious transactions can be detected quite well using these “simpler” features.
For the Elliptic data set, we see that the GNNs give a slightly worse performance in terms of AUC-ROC, except for the GAT, but mostly outperform the other models when looking at the more informative AUC-PR. The difference become apparent when looking at the threshold-dependent metrics in Tables 13E1E3 and E5. Precision, recall and F1-scores are remarkably higher among the observations with highest money laundering propensity.
One surprising result is the underperformance of the GIN method for the threshold-independent metrics. Two explanations are possible. The first is that the neural network in the aggregation step is not optimal, and should be tuned further. This was not part of this experimental set-up because this would add additional hyperparameter for the depth and size of the layers for each of the layers of the GNN, which would increase the number of possible value combinations massively. Another explanation is that the inclusion of neural network aggregation actually makes GIN more sensitive to the high class imbalance, because much more weights need to be trained. This might indicate that this model is less suited for anti-money laundering. It should be noted that good performance is observed for the precision, recall and F1-scores, indicating that GIN might still be a good model when only interested in reducing false positives in the top-percentiles of predictions.
For the IBM dataset, with AUC-ROC in Table 12, the picture is difference. Here, the network information coming from DeepWalk and node2vec give additional information and increase the performance of the model. However, when looking at the GNNs, the performance is bad. The performance of GAT has even dropped to an AUC-ROC of $0.27$, meaning that the inverse prediction actually holds more value. The threshold-dependent metrics show a similar picture, but in Tables 14E2E4 and E6 it seems that DeepWalk gives better performance in the top-percentiles.
Another anomalous observations is GraphSAGE, with perfect recall, but low precision. It seems that GraphSAGE is finding all money laundering cases, but at the cost of having a high fall positive rate. This might be caused by GraphSAGE giving a flat percentage for each observation. We are working with relative cut-offs, so having the cut-off equal to the top-$1\%$, classifies all observations as money laundering.
One explanation for the poor performance of GNNs might be related to the network structure, as presented in Section 4.2. Given that the average path length is small, and the nodes are very close to each other in the network, the GNN embedding might suffer from over-smoothing [79]. This can result in very similar embeddings for the different nodes with insufficient discriminatory power.
Moreover, a deeper analysis reveals that the training loss, when training on the train and validation set, is highly unstable. Depending on the epoch and specific weight initialization, the results vary widely, with the models sometimes just predicting the majority class. For this specific data set, GNNs seem to face instability due to the extreme class imbalance.
The supervised learning experiments on both data sets demonstrate that network analytics can provide additional insights for anti-money laundering. However, depending on the network topology and the extent of the class imbalance, GNNs might be too unstable. Additional experiments are needed to quantify the sensitivity of GNNs to the imbalance, but this is outside the scope of this research.
5.2 Unsupervised Learning
The results of the unsupervised methods using isolation forests are given in Tables 15. None of the results on the Elliptic data set are particularly good, with AUC-ROC values below 0.5 indicating that the inverse of the scores would perform better as a classifier. For completeness, we also include the threshold-dependent metric in Table 16, where most models fail to detect money laundering cases.
Further analysis is done to understand why the results for unsupervised learning might be this bad. Figure 6 shows the anomaly scores coming from the isolation forest trained on the intrinsic features for the Elliptic data set. Here, a low (very negative) score signifies anomalies. All money laundering cases have scores near zero, within the bulk of the distribution, indicating that the isolation forest does not recognize these cases as anomalies.
Conversely, the results for IBM-AML are more in line with expectations. Network information captured by DeepWalk, results in the best performing model here. Comparing these results to those in Table 12, we see that DeepWalk even outperforms the supervised learning methods in terms of AUC-PR.
Figure 6: The anomaly scores for the legitimate and money laundering cases. More negative is more anomalous.
Table 15: Threshold-independent metrics: AUC-ROC and AUC-PR values over the different methods for the two data set, based on the test set. The standard deviation is also reported.
Unsupervised Elliptic IBM-AML
Methods AUC-ROC AUC-PR AUC-ROC AUC-PR
Intrinsic features $0.134\pm 0.007$ $0.058\pm 0.001$ $0.710\pm 0.015$ $0.017\pm 0.003$
Egonet features $0.217\pm 0.010$ $0.013\pm 0.000$ $0.700\pm 0.033$ $0.011\pm 0.005$
DeepWalk $0.182\pm 0.008$ $0.056\pm 0.001$ $\boldsymbol{0.744\pm 0.023}$ $\boldsymbol{0.024\pm 0.008}$
Node2vec $0.190\pm 0.008$ $0.056\pm 0.0005$ $0.632\pm 0.031$ $0.005\pm 0.001$
@@ 5.2 Unsupervised Learning
Elliptic Unsupervised
Methods Precision Recall F1-score
Intrinsic features $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
Egonet features $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
DeepWalk $0.005\pm 0.005$ $0.001\pm 0.001$ $0.001\pm 0.001$
Node2vec $0.003\pm 0.004$ $000.000\pm 0$ 0.000 $0.001\pm 0.001$
Table 17: Threshold-dependent metrics: Precision, recall and F1-score values for the top $1\%$ scores over the different methods for the IBM-AML data set, based on the test set. The standard deviation is also reported.
IBM-AML Unupervised
Methods Precision Recall F1-score
Intrinsic features $0.042\pm 0.005$ $0.156\pm 0.020$ $0.067\pm 0.008$
Egonet features $0.027\pm 0.013$ $0.098\pm 0.048$ $0.042\pm 0.021$
DeepWalk $0.055\pm 0.009$ $0.204\pm 0.034$ $0.087\pm 0.015$
Node2vec $0.007\pm 0.002$ $0.025\pm 0.008$ $0.010\pm 0.003$
The most plausible explanation in our opinion lies in the assumptions underlying isolation forest and the data generating process. Two specific assumptions by Liu et al. [82]:
  • Anomalies yield fewer partitions since anomalies occupy regions with lower density;
  • Instances with distinct attributes are separated earlier in the tree, contributing to shorter paths.
The Elliptic data set is a real-life data set, where criminals actively attempt to cover their tracks [10]. By doing so, they actually become suspicious again. They try to mimic the ‘average behaviour’, but no real transaction/person is average on all attributes. In contrast, the IBM-AML data set is synthetic, with patterns and behaviours deliberately introduced. It is likely that the more complex camouflage behaviour is not present in the latter. This insight highlights the importance of knowing the exact assumptions underlying the data sets used. Synthetic data sets might give overly optimistic results, since complex behaviour inherent to the problem at hand are not captured.
@@ 6 Conclusion
In this paper, we present the results of a systematic literature review of the literature on network analytics for anti-money laundering, and of an experimental evaluation of the state-of-the-art. The systematic literature review resulted in a selection of 97 papers. These papers were processed in two steps to achieve (1) a comprehensive understanding of the broader literature and (2) detailed insight into research that has been well received by the community. The comprehensive analysis classified all papers according to seven categories, each having different sub-characteristics. The detailed analysis was performed on the $10\%$ most-cited papers in scope, which were further categorized using the framework recently proposed by Bockel-Rickermann et al. [14]. Information was given on the different processing steps taken, the challenges faced or addressed in the papers, and the definition of nodes and edges in the networks.
@@
We concluded that a large part of the literature concerns methods that rely heavily on expert-knowledge by applying manual feature engineering and rule-based methods. Additionally, most methods find patterns that are indicative for money laundering by relying on centrality metrics and other manually engineered features. However, most methods are specifically tailored to solving just a small part of the money laundering puzzle. This might indicate a latent consensus that one should not aim at developing a single, overarching model for detecting money laundering. Instead, one may benefit from operating multiple and complementary models in parallel, each detecting different money laundering patterns. This preference appears to be driven, in part, by the lack of interpretability in more complex network representation methods.
Our analysis identified five critical gaps that future research must address:
  1. 1.
    Interpretability: Few studies have explored tools to explain the outputs of graph neural networks (GNNs), which is essential for regulatory and operational transparency.
  2. 2.
    Unsupervised Methods: Limited exploration of classic unsupervised methods restricts their potential to detect anomalies, a key requirement as money laundering tactics evolve.
  3. 3.
    Dynamic Network Analysis: Most approaches rely on static networks, with insufficient research into dynamic methods that could better capture the temporal nature of transactions.
  4. 4.
    Incomplete Labels: The inherent incompleteness of AML labels, due to undetected cases and resource constraints, highlights the need for specialized PU-learning techniques.
  5. 5.
    Prioritized Predictions: Limited resources necessitate actionable, prioritized predictions, which could be achieved through learning-to-rank methods.
Additionally, for regular fraud analytics, cost-sensitive learning is an established set of methods [14, 134, 62, 107], while none of the papers covered in our work use the cost of misclassification in the learning of their methods. Although estimation of misclassification costs in money laundering is much harder than in transaction/credit card fraud, we believe it is a branch of the literature that is worth-wile exploring more.
To extend upon the existing body of knowledge, we implemented an extensive experimental evaluation covering manual feature engineering, shallow and deep representation learning, both in a supervised and unsupervised setting. From the supervised learning results, we can conclude that network features bring additional predictive power to AML models. Depending on the setting, GNNs perform better than the manual features.
In the unsupervised learning experiments, we employed isolation forests and observed notable differences between the real-world Elliptic dataset and the synthetic IBM-AML dataset. Although on synthetic data, the methods performed well, the observations from the Elliptic data seemed to exhibit strong camouflaging tactics. The anomaly scores of the latter were so low that investigating the least suspicious transactions for money laundering became the better strategy. We conclude from this that care should be taken when testing methods on synthetic data, since it can give overly optimistic results. Future work should also investigate if this aberration still holds for other unsupervised learning methods.
@@
For network visualisation, future work should focus on quantifying the usefulness of visualisations in helping investigators detect new patterns and money laundering cases. This can be extended into a comparative study that looks at the strengths and weaknesses of different visualisations. This requires qualitative empirical work studying how people use the visualisations, not just how it is intended to be used.
There are also some limitations in our experimental set-up that can be addressed in future work. First, conclusions are drawn based on only two data sets. To generalise our conclusions, additional tests need to be performed on a broader range of data sets. This can be done by including more open-source data sets when these become available, and rely on practitioners to use our code to test the methods and publish results on their proprietary data. Second, the decoders were always based on neural networks. The application and impact analysis of applying different machine learning methods as down-stream classifiers is another interesting area of study to extend the experiments presented here. Third, the hyperparameter tuning process was limited. While Optuna provided an efficient method for hyperparameter search, the 50 to 100 tuning rounds used in this sdy may be insufficient. The choices in this paper were made to keep everything tractable and feasible given the limited resources available. Future work could expand the parameter ranges and number of tuning rounds using more computing power.
@@ Acknowledgments
This work was supported by the Research Foundation – Flanders (FWO research project 1SHEN24N) and by the BNP Paribas Fortis Chair in Fraud Analytics. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government.
@@ Appendix A Literature Overview
Table A1: Literature classification. Classification of the papers according to different categories: publication data, learning method, evaluation metric, objective, and data.
Publication Data Evaluation Metric Objective Data Network
@@
Reference Year Review Paper Bitcoin/Crypto Accuracy Precision Recall F1 TPR FPR AUC-ROC AUC-PR Time Client Classification Transaction Classification Community Detection Flow/Chain Detection Proprietary Open-Source Synthetic Inverse/undirected network Multiple networks
Dreżewski et al. [33] 2015
Ngai et al. [100] 2011
Malm and Bichler [88] 2013
Fronzetti Colladon and Remondi [41] 2017
Zhou et al. [156] 2017
Didimo et al. [29] 2011
McGinn et al. [89] 2016
Senator et al. [116] 1995
Wu et al. [142] 2021
Gao and Ye [42] 2007
Joaristi et al. [69] 2018
Irwin and Turner [64] 2018
Michalak and Korczak [94] 2011
Wu et al. [143] 2021
Li et al. [80] 2020
Phetsouvanh et al. [106] 2018
Zhou et al. [159] 2017
Didimo et al. [31] 2019
Joaristi et al. [70] 2019
Camino et al. [19] 2017
Velasco et al. [136] 2021
Nandhini and Das [98] 2016
Zhdanova et al. [155] 2014
Imanpour et al. [63] 2019
Turner et al. [128] 2019
Ovelgönne et al. [102] 2012
Magalingam et al. [85] 2015
Helmy et al. [59] 2016
Xueshuo et al. [150] 2021
Plaksiy et al. [108] 2018
Garcia-Bedoya et al. [44] 2021
Mehmet and Wijesekera [91] 2013
Xia et al. [145] 2021
Kute et al. [75] 2021
Mahootiha et al. [87] 2021
Shaikh et al. [118] 2021
Turner et al. [129] 2020
Ricci [112] 2020
Bahulkar et al. [11] 2018
Chmielewski and Stkapor [24] 2017
Zhou et al. [157] 2022
Gasanova et al. [46] 2018
Magomedov et al. [86] 2018
Tharani et al. [125] 2021
Jin and Qu [68] 2018
Mehmet et al. [92] 2014
Xiao et al. [146] 2018
Mohan et al. [96] 2022
Alarab et al. [4] 2020
Shen et al. [119] 2021
Prado-Romero and Gago-Alonso [109] 2017
Zhang et al. [154] 2021
Garcia-Bernardo et al. [45] 2022
He et al. [58] 2021
Micale et al. [93] 2019
Bhalerao et al. [13] 2019
Omar et al. [101] 2014
Winiecki et al. [141] 2020
Starnini et al. [121] 2021
Hamid [54] 2017
Gulino et al. [52] 2021
Adriaens et al. [1] 2019
Chang et al. [22] 2007
Kurshan and Shen [74] 2020
Jin et al. [66] 2022
Bellomarini et al. [12] 2021
Sousa Lima et al. [120] 2022
Çağlayan and Bahti̇yar [17] 2022
Jin et al. [67] 2022
Kurniawan et al. [73] 2015
Sun et al. [122] 2022
Gerbrands et al. [47] 2022
Dumitrescu et al. [34] 2022
Yu et al. [153] 2022
García and Mateos [43] 2021
Wang et al. [138] 2022
Gong et al. [48] 2022
Li et al. [81] 2022
Li et al. [78] 2022
Assumpção et al. [8] 2022
Semenov et al. [115] 2017
Didimo et al. [30] 2012
Cardoso et al. [20] 2022
Day [27] 2021
Alarab et al. [5] 2020
Badalov et al. [9] 2018
Korczak and Łuszczyk [72] 2011
Lokanan [84] 2022
Rathore et al. [111] 2022
Liu et al. [83] 2012
Xiong and Zhong [148] 2020
Shaikh and Nazir [117] 2018
Yang et al. [151] 2022
Xiong and Xiao [147] 2021
Mehmet and Wijesekera [90] 2013
Cheong and Si [23] 2010
Han et al. [56] 2021
Table A2: Literature classification. Classification of the papers according to the different method building blocks.
Publication Data Method Type Modelling Method
Reference Year Bitcoin/Crypto Supervised Unsupervised Semi-Supervised Mixed Visualisation Rule-Based Manual Features Walk-based Shallow Repr. Deep Repr. Correlation-Based Tree-Based SVM-Based Neural Networks Anomaly Detection Clustering Logistic Regression
Dreżewski et al. [33] 2015
Malm and Bichler [88] 2013
Fronzetti Colladon and Remondi [41] 2017
Zhou et al. [156] 2017
Didimo et al. [29] 2011
McGinn et al. [89] 2016
Senator et al. [116] 1995
Wu et al. [142] 2021
Joaristi et al. [69] 2018
Michalak and Korczak [94] 2011
Wu et al. [143] 2021
Li et al. [80] 2020
Phetsouvanh et al. [106] 2018
Zhou et al. [159] 2017
Didimo et al. [31] 2019
Joaristi et al. [70] 2019
Camino et al. [19] 2017
Velasco et al. [136] 2021
Nandhini and Das [98] 2016
Zhdanova et al. [155] 2014
Imanpour et al. [63] 2019
Turner et al. [128] 2019
Ovelgönne et al. [102] 2012
Magalingam et al. [85] 2015
Helmy et al. [59] 2016
Xueshuo et al. [150] 2021
Plaksiy et al. [108] 2018
Garcia-Bedoya et al. [44] 2021
Mehmet and Wijesekera [91] 2013
Xia et al. [145] 2021
Mahootiha et al. [87] 2021
Shaikh et al. [118] 2021
Turner et al. [129] 2020
Ricci [112] 2020
Bahulkar et al. [11] 2018
Chmielewski and Stkapor [24] 2017
Zhou et al. [157] 2022
Gasanova et al. [46] 2018
Magomedov et al. [86] 2018
Tharani et al. [125] 2021
Jin and Qu [68] 2018
Mehmet et al. [92] 2014
Xiao et al. [146] 2018
Mohan et al. [96] 2022
Shen et al. [119] 2021
Prado-Romero and Gago-Alonso [109] 2017
Zhang et al. [154] 2021
Garcia-Bernardo et al. [45] 2022
He et al. [58] 2021
Micale et al. [93] 2019
Bhalerao et al. [13] 2019
Omar et al. [101] 2014
Winiecki et al. [141] 2020
Starnini et al. [121] 2021
Hamid [54] 2017
Gulino et al. [52] 2021
Adriaens et al. [1] 2019
Chang et al. [22] 2007
Jin et al. [66] 2022
Bellomarini et al. [12] 2021
Sousa Lima et al. [120] 2022
Çağlayan and Bahti̇yar [17] 2022
Jin et al. [67] 2022
Kurniawan et al. [73] 2015
Sun et al. [122] 2022
Dumitrescu et al. [34] 2022
Yu et al. [153] 2022
García and Mateos [43] 2021
Wang et al. [138] 2022
Gong et al. [48] 2022
Li et al. [81] 2022
Li et al. [78] 2022
Assumpção et al. [8] 2022
Didimo et al. [30] 2012
Cardoso et al. [20] 2022
Alarab et al. [5] 2020
Badalov et al. [9] 2018
Korczak and Łuszczyk [72] 2011
Rathore et al. [111] 2022
Liu et al. [83] 2012
Xiong and Zhong [148] 2020
Shaikh and Nazir [117] 2018
Yang et al. [151] 2022
Xiong and Xiao [147] 2021
Mehmet and Wijesekera [90] 2013
Cheong and Si [23] 2010